2. Based on the following information answer questions Holy products corporations have the following capital structure, which it considers optimal: Bonds, 7% (at par)      Br 300,000   

Preferred stock, Br.5     240,000

Common stock                        360,000 Retained earnings                      300,000   Additional Information: Dividends on common stock are currently Br 3 per share and are expected to grow at constant rate of 6%. Market price of common stock is Br 40 and the preferred stock is selling at Br50. Flotation cost on new issues of common stock is 10%. A. Based on the above information, what would be the cost of the bond? B. What would be the cost of common stock for Holy Products Corporation?


1
Expert's answer
2022-01-14T13:51:18-0500

Answer (A):

Cost of the bond = 7%


Answer (B):

The cost of common stock (Ke) for Holy Products Corporation is:


"Ke=\frac{D_1}{P_0(1-F)}+g"


"Ke=\frac{3\times(1+0.06)}{40\times(1-0.10)}+0.06"


"Ke=\frac{3.18}{36}+0.06"


"Ke=0.1483\approx14.83\%"